value. 1000 points 1 out of 2 attempts Enter your answer in the provided box. A 24.90 g sample of an alloy at 93.00°c is placed into 50.0 g of water at 22.00°c in an insulated coffee-cup calorimeter with a heat capacity of 9.20 JIK. If the final temperature of the system is 31.10 C, what is the specific heat capacity of the alloy? References eBook & Resources Multipart Answer Chapter 6 Thermochemistry Energy Flow and Chemicol Chang
Answer #1

Heat lost by the alloy= Heat gained by the calorimeter

malloycalloydeltaTalloy = mcalorieccalorideltaTcalorie

24.90g*calloy*(93-31.10)C = 50g*9.20J/K*(304.1-295)K

24.90g*61.9C*calloy =50g*9.1K*9.20J/K

1541.31Cg*calloy=4186Jg

calloy = 2.72J/C=2.72J/C /24.90g = 0.11J/Cg